### v1

# An interpretable family of projected normal distributions and a related copula model for Bayesian analysis of hypertoroidal data

Abstract
This paper introduces two families of probability distributions for Bayesian analysis of hypertoroidal data. The first family consists of symmetric distributions derived from the projection of multivariate normal distributions under specific parameter constraints. This family is closed under marginalization and hence any marginal distribution belongs to a lower-dimensional case of the same family. In particular the univariate marginal of the family is the unimodal case of the projected normal distribution on the circle. The second family is a flexible extension of the copula case of the first family, which can accommodate any univariate marginal distributions. Unlike existing models derived via projection, both families have the common advantage that their parameters possess a clear and intuitive interpretation. In addition, Markov Chain Monte Carlo algorithms are presented for Bayesian estimation of both families and a simulation study is used to demonstrate their performance. As a real data example, a meteorological data set is analyzed.
Keywords: Circular data; Circular dependence; Interpretable model.
1 Introduction
Data recorded as angles frequently appear in various academic fields. Such examples include wave directions in environmental science, directions of animal movement in biology, dihedral angles in protein backbone structures in bioinformatics, and times of crimes in political science. In practice, angular observations are often jointly observed with other angular observations, forming what is referred to as hypertoroidal data. For example, conformational angles in protein structure consist of two dihedral angles and one torsion angle of the side chain, and wave directions are typically observed at multiple locations in the sea. It is well-known that ordinary statistical methods for linear data cannot be directly applied to such hypertoroidal data and new statistical methods are needed for their analysis. See, for example, Mardia and Jupp, (1999) and Ley and Verdebout, (2017) for the general review of statistics of hypertoroidal data.
A major problem in analyzing general-dimensional toroidal data is the lack of tractable and interpretable probability distributions that can serve as a foundation for parametric statistical analysis. However, if the analysis is restricted to 2- or 3-dimensional toroidal cases, there are a few distributions that are both tractable and easily interpreted. The well-known families of distributions for the 2-dimensional case are the bivariate von Mises distributions (Mardia,, 1975; Singh et al.,, 2002; Mardia et al.,, 2007; Kent et al.,, 2008; Mardia,, 2010) and copula-based models (Wehrly and Johnson,, 1980; Shieh and Johnson,, 2005; Shieh et al.,, 2011; Jones et al.,, 2015; Kato and Pewsey,, 2015). A 3-dimensional extension of Wehrly and Johnson, (1980) using the wrapped Cauchy-type copula density was recently proposed by Kato et al., (2024). Several distributions on the general-dimensional torus have been proposed in the literature. Mardia et al., (2008) and Mardia and Voss, (2014) considered a multivariate extension of the von Mises distribution, and Navarro et al., (2017) discussed its generalization. Using transformations of the multivariate normal distribution, the multivariate wrapped distribution (Jona-Lasinio et al.,, 2012), the projected normal (PN) distribution (Mastrantonio,, 2018), and the inverse stereographic normal distribution (Selvitella,, 2019) have been proposed. A tree-structured model on the general-dimensional torus was proposed by Leguey et al., (2016) in machine learning. Models on the hypertorus using other approaches include the models of Fernández-Durán and Gregorio-Domínguez, (2014) and Kim et al., (2016). Note also that, apart from the models of Leguey et al., (2016), Navarro et al., (2017) and Mastrantonio, (2018), Bayesian analysis of these multivariate distributions has not been sufficiently explored. However, these existing distributions for general-dimensional toroidal data have some limitations, such as unclear parameter interpretations, multimodality, estimations, or limited dependence structure.
In this paper we propose two families of probability distributions for the Bayesian analysis of general-dimensional toroidal data. The first family, the toroidal projected normal (TPN) distribution, is derived by projecting multivariate normal distributions under specific parameter constraints. The TPN model inherits several tractable properties of the multivariate normal distribution. For example, the TPN distribution is symmetric on the torus and, in addition, the family of TPN distributions is closed under marginalization. In particular, the univariate marginal of the TPN distribution is the symmetric and unimodal case of the projected normal (PN) distribution (see, e.g., Mardia and Jupp,, 1999, Equation (3.5.49)). The second family proposed in the paper is an extension of the copula case of the TPN (CTPN) family. The CTPN is a flexible family of distributions with arbitrary univariate marginal distributions and it can be thought as a multivariate extension of the bivariate model of Wehrly and Johnson, (1980). The modes of the joint CTPN density can be expressed in a simple form under certain parameter conditions. Even though a closed-form expression for the copula density is not available in the general case, we derive it in closed form for the bivariate case. Given that the TPN, and hence also its copula, is closed under marginalization, this shows that regardless of the data dimension, the two-dimensional copula, which we know in closed form, represents the exact bivariate dependence we are assuming. For the TPN family, the parameters have clear and intuitive interpretations: each marginal mean direction and concentration is ruled by a specific parameter, and we have a parameter to model each correlation between pairs of variables. For the CTPN families, the mean directions and concentrations of the univariate marginals are determined by the parameters of the marginal distributions, while the dependence among circular variables is governed by the copula density. Similar to the TPN family, the copula parameters that rule the dependence, have clear and intuitive interpretations. Furthermore, due to their connection with multivariate normal distributions, both proposed families are well-suited for Bayesian estimation. We define a new prior distribution for the covariance matrix of the underlying normal distribution, needed for the constrains that must be imposed to obtain interpretable models, and we develop Markov chain Monte Carlo (MCMC) algorithms for Bayesian estimation of the proposed families. With a simulation study, we show that the MCMC algorithm is able to retrieve the values of the parameters used to simulate the data for both the TPN and a copula case. The proposed distributions are estimated on real meteorological data, where time series of four wind direction and two wave direction measurements are recorded.
It should be noted that the TPN is a submodel of the circular-linear distribution of Mastrantonio, (2018). The hypertoroidal submodel of Mastrantonio, (2018) offers greater flexibility than our TPN distribution, as it allows for multimodality and non-symmetry. However, as with other projected distributions (see Watson,, 1983, Section 3.6; Wang and Gelfand,, 2014; Mastrantonio et al.,, 2016), this flexibility comes at the cost of parameter interpretability. On the other hand, a tractable, interpretable, and flexible distribution can serve as a solid foundation for parametric statistical analysis. Moreover, only by constraining the parameters of the PN, it is possible to obtain a copula model with a closed-form marginal bivariate density.
The paper is organized as follows: in Section 2, we define the family of TPN distributions and provide its basic properties; in Section 3, we propose the copula version of the TPN distribution and its extension with arbitrary univariate marginal distributions; Section 4 discusse the circular correlation of our proposals; Section 5 presents the MCMC algorithm for the TPN distribution and its related copula model; in Section 6, we analyze a real dataset using our proposals; Section 7 concludes the paper. The proofs of the theoretical claims and the simulation experiments are provided in the Online Material, available on the journal website. The code used to produce the results of the paper is available at https://github.com/GianlucaMastrantonio/toroidal_projected_normal.
2 A Family of Projected Normal Distributions on the Hypertorus
2.1 Idea behind the derivation of the proposed family
Before introducing the proposed family and its properties, we first explain the idea behind its derivation using the univariate and bivariate settings. In the univariate setting, a circular variable $\Theta$ is said to follow a projected normal (PN) distribution, denoted by $\mathrm{PN}((\eta_{c},\eta_{s})^{\top},\bm{\Psi})$, if a bivariate random vector $(X_{c},X_{s})^{\top}$ follows a bivariate normal distribution $N_{2}((\eta_{c},\eta_{s})^{\top},\bm{\Psi})$ and $\Theta$ is defined through the transformation:
$X_{c}=R\cos\Theta,\quad X_{s}=R\sin\Theta.$ (1)
The angle, which is the circular variable of interest, can be expressed as
$\Theta=\text{atan}^{*}\left(\frac{X_{s}}{X_{c}}\right),$ (2)
where $\text{atan}^{*}(y/x)$ is the modified inverse tangent function defined as an angle in $\pi,\pi)$ whose sine and cosine are given by $y/\{x^{2}+y^{2}\}^{1/2}$ and $x/\{x^{2}+y^{2}\}^{1/2}$, respectively. It is known that when $\bm{\Psi}=\mathbf{I}_{2}$ (Mardia and Jupp,, 1999, Equation (3.5.49)), the PN is symmetric around the mean direction, which is given by
$\mu=\text{atan}^{*}\left(\frac{\eta_{s}}{\eta_{c}}\right),$ (3)
and the circular concentration
$\kappa=\sqrt{\eta_{c}^{2}+\eta_{s}^{2}}$ (4)
is a function of the mean vector $(\eta_{c},\eta_{s})^{\top}$. The starting point of our proposal, which to the best of our knowledge has never been pointed out before, is that, under the assumption $\bm{\Psi}=\mathbf{I}_{2}$, we can construct a PN-distributed variable by directly using $(\mu,\kappa)$ as parameters. This is possible if we assume that $(X_{c},X_{s})^{\top}\sim N((\kappa,0)^{\top},\mathbf{I}_{2})$, with
$X_{c}=R\cos(\Theta-\mu),\quad X_{s}=R\sin(\Theta-\mu).$ (5)
We can now move to the bivariate setting. If we have two circular variables $\Theta_{1}$ and $\Theta_{2}$, which are derived from the linear variables
$\bm{X}_{c}=\begin{pmatrix}X_{c,1}\\ X_{c,2}\end{pmatrix}=\begin{pmatrix}R_{1}\cos(\Theta_{1}-\mu_{1})\\ R_{2}\cos(\Theta_{2}-\mu_{2})\end{pmatrix}\quad\mbox{and}\quad\bm{X}_{s}=\begin{pmatrix}X_{s,1}\\ X_{s,2}\end{pmatrix}=\begin{pmatrix}R_{1}\sin(\Theta_{1}-\mu_{1})\\ R_{2}\sin(\Theta_{2}-\mu_{2})\end{pmatrix}$, (6)
an easy way to introduce dependence between the circular variables while keeping the marginal distribution invariant (see Theorem 2 below for a formal proof) is to let $\bm{X}_{c}\perp\bm{X}_{s}$ and assume
$\bm{X}_{c}\sim N_{d}(\bm{\kappa},\bm{\Sigma}_{c}),\quad\bm{X}_{s}\sim N_{d}(\bm{0},\bm{\Sigma}_{s}),$ (7)
where $\bm{\kappa}\in(\mathbb{R}^{+})^{2}$, and both $\bm{\Sigma}_{c}$ and $\bm{\Sigma}_{s}$ are correlation matrices. Depending on the values of $\rho_{c}=[\bm{\Sigma}_{c}]_{1,2}$ and $\rho_{s}=[\bm{\Sigma}_{s}]_{1,2}$, we may observe different types of dependence. For interpretability purposes and parsimony of parameters, we want to have a single parameter $\rho$ such that, for positive values, the two circular variables tend to rotate in the same direction, while for negative values, they tend to rotate in opposite directions.
From Wang and Gelfand, (2013) we know that $E(\sin(\Theta_{1}-\mu_{1}))=E(\sin(\Theta_{2}-\mu_{2}))=0$. It is evident that $X_{s,j}$ and $\sin(\Theta_{1}-\mu_{1})$ have the same sign and, moreover, the larger the value of $X_{s,j}$, the larger $\sin(\Theta_{1}-\mu_{1})$ is. Hence, a positive $\rho_{s}$ will induce a positive correlation between the two sine transformations while a negative one will have the opposite effect. To induce the correct correlation between the two circular variables, we also need to define the correlation between the two $X_{c,j}$, which induces a dependence between the cosines. In this case, regardless of whether the relationship between $\sin(\Theta_{1}-\mu_{1})$ and $\sin(\Theta_{2}-\mu_{2})$ is proportional or inversely proportional, as $\Theta_{j}-\mu_{j}\in[-\pi,\pi)$ decreases within $\pi,0]$ or increases within $0,\pi)$, $\cos(\Theta_{j}-\mu_{j})$ decreases, as does $X_{c,j}$. This suggests that, for the sign of $\rho_{c}$ to reflect the relationship between $\Theta_{1}$ and $\Theta_{2}$, the parameter $\rho_{c}$ must always be positive. For this reason, we assume $\bm{\Sigma}_{s}]_{1,2}=\rho$ and $\bm{\Sigma}_{c}]_{1,2}=|\rho|$, which is the basic idea of our proposal. A generalization to arbitrary dimension is proposed below.
2.2 Definition and basic properties
Definition 1.
Let $\bm{X}_{c}$ and $\bm{X}_{s}$ be two independent $d$-dimensional random vectors distributed as $N_{d}(\bm{\kappa},\bm{\Sigma}_{c})$ and $N_{d}(\bm{0},\bm{\Sigma})$, respectively, where $\bm{\kappa}=(\kappa_{1},\ldots,\kappa_{d})^{\top}$, $\kappa_{1},\ldots,\kappa_{d}\geq 0$, and $\bm{\Sigma}_{c}$ and $\bm{\Sigma}$ are correlation matrices with elements $\bm{\Sigma}]_{j,k}=\rho_{jk}$ and $\bm{\Sigma}_{c}]_{j,k}=|\rho_{jk}|$. Transform $\bm{X}_{c}$ and $\bm{X}_{s}$ into
$\bm{X}_{c}=\begin{pmatrix}R_{1}\cos(\Theta_{1}-\mu_{1})\\ R_{2}\cos(\Theta_{2}-\mu_{2})\\ \vdots\\ R_{d}\cos(\Theta_{d}-\mu_{d})\end{pmatrix}\quad\mbox{and}\quad\bm{X}_{s}=\begin{pmatrix}R_{1}\sin(\Theta_{1}-\mu_{1})\\ R_{2}\sin(\Theta_{2}-\mu_{2})\\ \vdots\\ R_{d}\sin(\Theta_{d}-\mu_{d})\end{pmatrix}$, (8)
with $R_{j}>0$, $\Theta_{j}\in[-\pi,\pi)$, and $\pi\leq\mu_{1},\ldots,\mu_{d}<\pi$. Then the random vector $(\Theta_{1},\ldots,\Theta_{d})$ consists of $d$-dimensional circular variables living on the hypertorus $\pi,\pi)^{d}$, and we refer to their joint distribution as the (hyper-)toroidal projected normal (TPN) distribution.
For convenience, we write $(\Theta_{1},\ldots,\Theta_{d})\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$, with $\bm{\mu}=(\mu_{1},\ldots,\mu_{d})^{\top}$, if the random vector $(\Theta_{1},\ldots,\Theta_{d})$ follows the distribution in Definition 1. It should be noted that both $\bm{\Sigma}$ and $\bm{\Sigma}_{c}$ are positive definite correlation matrices. However, not every matrix $\bm{\Sigma}$ produces a matrix $\bm{\Sigma}_{c}$ that is positive definite. This poses challenges for the estimation of the covariance matrices since we need to restrict the space where $\bm{\Sigma}$ is defined. We will tackle this issue with an appropriate prior distribution in Section 5, where the MCMC algorithm is presented.
From the definition is easy to prove the following Theorem:
Theorem 1.
Let $\bm{\Theta}\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$. Then the distribution of $\bm{\Theta}$ is symmetric about $\bm{\mu}$, namely, $\bm{\Theta}-\bm{\mu}\stackrel{{\scriptstyle d}}{{=}}\bm{\mu}-\bm{\Theta}$.
The symmetry of our model leads to clear interpretation of the parameters, which include not only $\bm{\mu}$ itself but also $\bm{\kappa}$ and $\bm{\Sigma}$; see Theorem 3 and Figures 1 and 2 below.
Theorem 2.
Let $(\Theta_{1},\ldots,\Theta_{d})\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$. Suppose $\bm{\Theta}_{1}=(\Theta_{1},\ldots,\Theta_{d_{1}})$ and $\bm{\Theta}_{2}=(\Theta_{d_{1}+1},\ldots,\Theta_{d})$. Then
$\bm{\Theta}_{1}\sim TPN_{d_{1}}(\bm{\mu}_{1},\bm{\kappa}_{1},\bm{\Sigma}_{11})\quad\text{and}\quad\bm{\Theta}_{2}\sim TPN_{d_{2}}(\bm{\mu}_{2},\bm{\kappa}_{2},\bm{\Sigma}_{22}),$
where $\bm{\mu}_{1}=(\mu_{1},\ldots,\mu_{d_{1}})^{\top}$, $\bm{\mu}_{2}=(\mu_{d_{1}+1},\ldots,\mu_{d})^{\top}$, $\bm{\kappa}_{1}=(\kappa_{1},\ldots,\kappa_{d_{1}})^{\top}$, $\bm{\kappa}_{2}=(\kappa_{d_{1}+1},\ldots,$ $\kappa_{d})^{\top}$, $d_{1}+d_{2}=d$, and $\bm{\Sigma}_{11}$ and $\bm{\Sigma}_{22}$ are $d_{1}\times d_{1}$ and $d_{2}\times d_{2}$ sub-matrices of $\bm{\Sigma}$, respectively, defined by
$\bm{\Sigma}=\begin{pmatrix}\bm{\Sigma}_{11}&\bm{\Sigma}_{12}\\ \bm{\Sigma}_{12}^{\top}&\bm{\Sigma}_{22}\end{pmatrix}$,
with $\bm{\Sigma}_{12}$ being a $d_{1}\times d_{2}$ matrix.
Note that any permutation of the random vector $(\Theta_{1},\ldots,\Theta_{d})$ also has the TPN distribution due to the symmetric form of the distribution and is hence closed under marginalization.
Theorem 3.
Let $(\Theta_{1},\ldots,\Theta_{d})\sim TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma})$. Then the marginal distribution of $\Theta_{j}$ $(j=1,\ldots,d)$ is a symmetric and unimodal case of the PN distribution, with density
$f(\theta_{j})=\frac{1}{\sqrt{2\pi}}\phi(\kappa_{j})+\kappa_{j}\cos(\theta_{j}-\mu_{j})\phi(\kappa_{j}\sin(\theta_{j}-\mu_{j}))\Phi(\kappa_{j}\cos(\theta_{j}-\mu_{j})),$ (9)
where $\phi$ and $\Phi$ denote the probability density function and cumulative distribution function of the standard normal distribution, respectively.
The proof for Theorem 3 is straightforward and follows directly from Mardia and Jupp, (1999, Section 3.5.6): it is therefore omitted.
From Theorems 13, it follows that our model can be considered a toroidal extension of the symmetric and unimodal PN distribution on the circle. Theorem 2 also gives the distribution of any possible combination of two circular variables, which are marginally TPN distributed. Then, given the discussion in Section 2.1, we can clearly see that $\mu_{j}$ is the mean direction of $\Theta_{j}$, the concentration of $\Theta_{j}$ depends only on $\kappa_{j}$, while the magnitude and sign of $\rho_{jk}$ measure the dependence between $\Theta_{j}$ and $\Theta_{k}$. While the TPN distribution is a special case of the poly-cylindrical distribution of Mastrantonio, (2018), our model, in comparison with Mastrantonio, (2018), has the advantage that each parameter can be clearly interpreted. Plots of the univariate and bivariate densities of the TPN are shown in Figures 1 and 2. The densities in Figure 1 are computed using (9), while those in Figure 2 are computed via numerical integration of the expression (12) given later.
(a)                            (b)
Figure 1: Plot of the density (9) for: (a) $\kappa_{1}=0.49$ with $\mu_{1}=0$ (solid), $\mu_{1}=\pi/3$ (dashed), $\mu_{1}=2\pi/3$ (dotted), and $\mu_{1}=-2\pi/3$ (dot-dashed); and (b) $\mu_{1}=0$ with $\kappa_{1}=0$ (dot-dashed), $\kappa_{1}=0.49$ (dotted), $\kappa_{1}=1.1$ (dashed), and $\kappa_{1}=2.45$ (solid).
$\kappa_{1}=0$          $\kappa_{1}=0.49$         $\kappa_{1}=1.1$         $\kappa_{1}=2.45$
$\rho_{12}=0$    Refer to caption Refer to caption Refer to caption Refer to caption
$\rho_{12}=0.3$    Refer to caption Refer to caption Refer to caption Refer to caption
$\rho_{12}=0.6$    Refer to caption Refer to caption Refer to caption Refer to caption
$\rho_{12}=0.9$Refer to caption Refer to caption Refer to caption Refer to caption
       Refer to caption Refer to caption Refer to caption Refer to caption
         (i)             (ii)            (iii)            (iv)
Figure 2: Contour plots of density (12) of $(\Theta_{1},\Theta_{2})$ for: (a) $\mu_{1}=\mu_{2}=0$, $\kappa_{2}=0.49$, with various combinations of $(\kappa_{1},\rho_{12})$; and (b) the same parameters as in the upper next frame of (a), but with the following adjustments: (i) $\mu_{1}=0$ is replaced by $\mu_{1}=\pi/2$, (ii) $\mu_{2}=0$ is replaced by $\mu_{2}=-\pi/2$, (iii) $\rho_{12}=0.9$ is replaced by $\rho_{12}=-0.9$, and (iv) $\kappa_{2}=0.49$ is replaced by $\kappa_{2}=2.45$.
Figure 1 shows the univariate marginal density (9) of $\Theta_{1}$ for selected values of the parameter. The four values of $\kappa_{1}$ used in Figure 1, namely, $(0,0.49,1.1,2.45)$, approximately correspond to mean resultant lengths $|E(e^{{i}\Theta_{1}})|$ of $(0,0.3,0.6,0.9)$, respectively. The mean resultant length is a measure of concentration for circular random variable, ranging from zero for uniformity to one for a point mass (see, e.g., Mardia and Jupp,, 1999, Section 3.4.2). As seen in Figure 1(a), the parameter $\mu_{1}$ represents the mode of the univariate density (9). The density (9) reaches its maximum value at $\theta_{1}=\mu_{1}$ and its minimum value at $\theta_{1}=\mu_{1}+\pi$. It is also clear that the density (9) is symmetric about $\theta_{1}=\mu_{1}$. Figure 1(b) shows how parameter $\kappa_{1}$ controls the concentration of the univariate marginal distribution. If $\kappa_{1}=0$, then the marginal distribution of $\Theta_{1}$ is uniform on the circle. The greater the value of $\kappa_{1}$, the greater the concentration of the marginal density of $\Theta_{1}$ around $\theta_{1}=\mu_{1}$. It can be shown that, as $\kappa_{1}\to\infty$, the marginal distribution of $\Theta_{1}$ converges to a point distribution with a singularity at $\theta_{1}=\mu_{1}$.
Figure 2 displays the bivariate marginal density of TPN of $(\Theta_{1},\Theta_{2})$, whose functional form will be given later in (12), for selected combinations of the parameters. In Figure 2(a), for fixed $\kappa_{1}$, it can be seen that the parameter $\rho_{12}$ controls the strength of dependence between $\Theta_{1}$ and $\Theta_{2}$. When $\rho_{12}=0$, $\Theta_{1}$ and $\Theta_{2}$ are independent. As $|\rho_{12}|$ increases, the strength of dependence between $\Theta_{1}$ and $\Theta_{2}$ also increases. The comparison between frame (b)(iii) and the frame directly above it in (a) shows that positive and negative values of $\rho_{12}$ imply positive and negative associations between the two variables, respectively. The role of the parameter $\kappa_{1}$ is evident from Figure 2(a) for fixed $\rho_{12}$, enhancing the interpretation of $\kappa_{1}$ for the univariate marginal given in Figure 1(b). When $\kappa_{1}=0$, the TPN is highly dispersed with respect to $\theta_{1}$. As $\kappa_{1}$ increases, the density becomes more concentrated around $\theta_{1}=\mu_{1}$. The parameter $\kappa_{2}$ plays an analogous role with respect to $\theta_{2}$; see frame (iv) of Figure 2(b). Finally, frames (i) and (ii) of Figure 2(b), in comparison with the frames directly above them, suggest that the joint density achieves its maximum at $(\theta_{1},\theta_{2})=(\mu_{1},\mu_{2})$.
2.3 Probability density function
Although the probability density function of the TPN cannot be written in closed form, here we show some new results that are also useful for the definition of the copula model based of the TPN. Let $\bm{R}=(R_{1},\ldots,R_{d})^{\top}$, then it is straightforward to see that the joint density of $(\bm{R},\bm{\Theta})$ can be expressed as
$f(\bm{r},\bm{\theta})$ $=f_{\bm{X}_{c}}(\bm{x}_{c})f_{\bm{X}_{s}}(\bm{x}_{s})\prod_{j=1}^{d}r_{j}$
$=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\exp\left[-\frac{1}{2}\left\{(\bm{x}_{c}-\bm{\kappa})^{\top}\bm{\Sigma}_{c}^{-1}(\bm{x}_{c}-\bm{\kappa})+\bm{x}_{s}^{\top}\bm{\Sigma}^{-1}\bm{x}_{s}\right\}\right]\prod_{j=1}^{d}r_{j},$ (10)
where $\bm{x}_{c}$ and $\bm{x}_{s}$ in (10) must be seen as functions of $(\bm{r},\bm{\theta})$, as defined in equation (8). It follows that the density of $\bm{\Theta}$ is
$\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}\left\{(\bm{x}_{c}-\bm{\kappa})^{\top}\bm{\Sigma}_{c}^{-1}(\bm{x}_{c}-\bm{\kappa})+\bm{x}_{s}^{\top}\bm{\Sigma}^{-1}\bm{x}_{s}\right\}\right]\prod_{j=1}^{d}r_{j}d\bm{r}.$ (11)
As is the case for the PN distribution on the hypersphere, it is difficult to obtain a closed form expression for $f(\bm{\theta})$ in the general case, but we will show that under our parametrization, it is possible to do it in some special case, see Section 3, where the copula model based on the TPN is introduced. Using a method similar to the one in Mastrantonio, (2018), this form of the density will be useful for the implementation of the MCMC algorithm for both the TPN and the copula case, see Section 5.
The density (11) has an alternative expression which is given below.
Theorem 4.
The density (11) has the alternative expression
$\begin{split}f(\bm{\theta})=&\ \frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\exp\left[-\frac{1}{2}\bm{\kappa}^{\top}\left(\mathbf{I}-\bm{\Sigma}_{c}^{-1}\mathbf{T}_{c}\mathbf{V}\mathbf{T}_{c}\right)\bm{\Sigma}_{c}^{-1}\bm{\kappa}\right]\\ &\times\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}(\bm{r}-\bm{\eta})^{\top}\mathbf{V}^{-1}(\bm{r}-\bm{\eta})\right]\prod_{j=1}^{d}r_{j}\,d\bm{r},\end{split}$ (12)
where $\bm{\eta}=\mathbf{V}\mathbf{T}_{c}\bm{\Sigma}_{c}^{-1}\bm{\kappa}$, $\mathbf{V}=(\mathbf{T}_{c}\bm{\Sigma}_{c}^{-1}\mathbf{T}_{c}+\mathbf{T}_{s}\bm{\Sigma}^{-1}\mathbf{T}_{s})^{-1}$, $\mathbf{T}_{c}=\mbox{\rm diag}(\cos(\theta_{1}-\mu_{1}),\ldots,\cos(\theta_{d}-\mu_{d}))$, and $\mathbf{T}_{s}=\mbox{\rm diag}(\sin(\theta_{1}-\mu_{1}),\ldots,\sin(\theta_{d}-\mu_{d}))$.
Since the exponential inside the integral of (12) corresponds to the kernel of a multivariate truncated normal distribution with parameters $\{\bm{\eta},\mathbf{V}\}$ for the variable $\bm{r}$, the integral is equal to
$E(R_{1}R_{2}\dots R_{d})\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}(\bm{r}-\bm{\eta})^{\top}\mathbf{V}^{-1}(\bm{r}-\bm{\eta})\right]d\bm{r},$ (13)
where $E(R_{1}R_{2}\dots R_{d})$ is the expectation with respect to the multivariate truncated normal distribution on the positive orthant, and the integral in (13) is the normalizing constant of the truncated normal density. Although $E(R_{1}R_{2}\dots R_{d})$ cannot be computed in closed form, numerical methods exist that allow to obtain its value, see Tallis, (1961), Lee, (1983) and Leppard and Tallis, (2018). The integral part is proportional to the value of the cumulative distribution function of the multivariate normal distribution $N_{d}(-\bm{\eta},\mathbf{V})$ at the origin, and can be computed straightforwardly using numerical methods.
As we can see from the following theorem, there are condition in which the matrix $\mathbf{V}$ in Theorem 4 can be greatly simplified:
Corollary 1.
Suppose that there exists $(\delta_{1},\ldots,\delta_{d})\in\{-1,1\}^{d}$ such that $\bm{\Sigma}_{c}=\mathbf{D}\bm{\Sigma}\mathbf{D}$, where $\mathbf{D}=\mbox{\rm diag}(\delta_{1},\ldots,\delta_{d})$. Then $\mathbf{V}$ is of the form
$\mathbf{V}=\begin{pmatrix}s_{11}&s_{12}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{2}\bar{\theta}_{2})&\cdots&s_{1d}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{d}\bar{\theta}_{d})\\ s_{12}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{2}\bar{\theta}_{2})&s_{22}&\cdots&s_{2d}\cos(\delta_{2}\bar{\theta}_{2}-\delta_{d}\bar{\theta}_{d})\\ \vdots&\vdots&\ddots&\vdots\\ s_{1d}\cos(\delta_{1}\bar{\theta}_{1}-\delta_{d}\bar{\theta}_{d})&s_{2d}\cos(\delta_{2}\bar{\theta}_{2}-\delta_{d}\bar{\theta}_{d})&\cdots&s_{dd}\end{pmatrix}^{-1},$
where $\bar{\theta}_{j}=\theta_{j}-\mu_{j}$ and $s_{jk}$ is the $(j,k)$ element of $\bm{\Sigma}_{c}^{-1}$.
Consider the bivariate marginal distribution of $(\Theta_{j},\Theta_{k})$ $(j\neq k)$. Denote the submatrix of $\bm{\Sigma}$ and $\bm{\Sigma}_{c}$ by
$\bm{\Sigma}^{(j,k)}=\begin{pmatrix}1&\rho_{jk}\\ \rho_{jk}&1\end{pmatrix},\quad\bm{\Sigma}_{c}^{(j,k)}=\begin{pmatrix}1&|\rho_{jk}|\\ |\rho_{jk}|&1\end{pmatrix}$.
Then there always exists $(\delta_{j},\delta_{k})$ such that $\bm{\Sigma}^{(j,k)}_{c}=\mathbf{D}\bm{\Sigma}^{(j,k)}\mathbf{D}$. In fact, by assuming $\delta_{j}=1$ and $\delta_{k}=\text{sign}(\rho_{jk})$, it is easy to see that $\bm{\Sigma}^{(j,k)}_{c}=\mathbf{D}\bm{\Sigma}^{(j,k)}\mathbf{D}$ holds for any $\rho_{jk}$, where $\mbox{sign}(\rho)=1$ if $\rho\geq 0$ and $\mbox{sign}(\rho)=-1$ if $\rho<0$. Thus Theorems 2 and 4 imply that the bivariate marginal density of $(\Theta_{j},\Theta_{k})$ can be expressed in the same form as (12) with $d=2$, and Corollary 1 suggests that the matrix $\mathbf{V}$ which corresponds to the bivariate case is given by
$\mathbf{V}$ $=\left\{\frac{1}{1-\rho_{jk}^{2}}\begin{pmatrix}1&\tau(\theta_{j},\theta_{k})\\ \tau(\theta_{j},\theta_{k})&1\end{pmatrix}\right\}^{-1}=\frac{1-\rho_{jk}^{2}}{1-\tau(\theta_{j},\theta_{k})^{2}}\begin{pmatrix}1&-\tau(\theta_{j},\theta_{k})\\ -\tau(\theta_{j},\theta_{k})&1\end{pmatrix}$, (14)
where
$\tau(\theta_{j},\theta_{k})=-|\rho_{jk}|\cos(\bar{\theta}_{j}-\mbox{sign}(\rho_{jk})\bar{\theta}_{k}).$ (15)
3 An Extension of the Copula Case of the TPN Distribution
3.1 The copula case of the TPN distribution
Let a random vector $\bm{\Theta}=(\Theta_{1},\ldots,\Theta_{d})^{\top}$ follow $TPN_{d}(\bm{\mu},\bm{0},\bm{\Sigma})$. Then, from (11) and (12), it follows that the density of $\bm{\Theta}$ reduces to
$\begin{split}c(\bm{\theta})&=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left[-\frac{1}{2}\left(\bm{x}_{c}^{\top}\bm{\Sigma}_{c}^{-1}\bm{x}_{c}+\bm{x}_{s}^{\top}\bm{\Sigma}^{-1}\bm{x}_{s}\right)\right]\prod_{j=1}^{d}r_{j}d\bm{r}\\ &=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\int_{(\mathbb{R}^{+})^{d}}\exp\left(-\frac{\bm{r}^{\top}\mathbf{V}^{-1}\bm{r}}{2}\right)\prod_{j=1}^{d}r_{j}\,d\bm{r}.\end{split}$ (16)
It should be noted that, in the second expression of the density (16), the variable $\bm{\theta}$ appears only inside the matrix $\mathbf{V}$. Then, under the assumption in Corollary 1, the density (16) is a function of $(\delta_{i}\bar{\theta}_{i}-\delta_{j}\bar{\theta}_{j})_{1\leq i,j\leq d}$.
It immediately follows from Theorem 3 that each univariate marginal of this submodel is the uniform distribution on the circle with density
$c_{1}(\theta_{k})=\frac{1}{2\pi}.$ (17)
This means that (16) is the density of a hypertoroidal copula model. Following standard practice in copula theory, we denote the density of the copula case of our family by $c$.
The modes of the copula case in (16) become clear under certain conditions.
Theorem 5.
Suppose that there exists $(\delta_{1},\ldots,\delta_{d})\in\{-1,1\}^{d}$ such that $\bm{\Sigma}_{c}=\mathbf{D}\bm{\Sigma}\mathbf{D}$, where $\mathbf{D}=\mbox{\rm diag}(\delta_{1},\ldots,\delta_{d})$. Let $s_{jk}<0$ for $1\leq j,k\leq d,\ j\neq k$, where $s_{jk}$ denotes the $(j,k)$ element of $\bm{\Sigma}_{c}^{-1}$. Then the copula density (16) takes its maximum values if and only if $\delta_{1}(\theta_{1}-\mu_{1})=\delta_{2}(\theta_{2}-\mu_{2})=\cdots=\delta_{d}(\theta_{d}-\mu_{k})$.
Note that the assumption of this theorem always holds when $d=2$. Although this is not the case for $d\geq 3$ in general, our simulation results suggest that the condition of the theorem often holds for $d=3$ for randomly generated $\bm{\Sigma}$ matrices with positive off-diagonal elements.
Interestingly, the bivariate marginal densities of the copula family have simple functional form.
Theorem 6.
The bivariate marginal density of $(\Theta_{j},\Theta_{k})$ has the closed-form expression:
$\begin{split}c_{2}(\theta_{j},\theta_{k})=&\ \frac{1}{(2\pi)^{2}}\frac{1-\rho^{2}}{\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{3/2}}\\ &\times\left\{\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}-\tau(\theta_{j},\theta_{k})\,\text{atan}^{*}\left(\frac{\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}}{\tau(\theta_{j},\theta_{k})}\right)\right\},\end{split}$ (18)
where $\tau(\theta_{i},\theta_{j})$ is given in (15) and $\text{atan}^{*}$ is defined in Section 2.1.
This theorem implies that the distribution (16) can be considered a multivariate extension of the bivariate copula model of Wehrly and Johnson, (1980), which is defined by a density that depends only on $\theta_{1}-\theta_{2}$ or $\theta_{1}+\theta_{2}$. The model of Wehrly and Johnson, (1980) has played an important role in directional statistics as a general model for bivariate circular data; see Jones et al., (2015) for detailed properties of their model. Note that the closed-form expression (18) of the bivariate marginal comes from the fact that the conditions of Corollary 1 are always met in this case. It is also possible to avoid the use of $\text{atan}^{*}$ and adopt the ordinary arctangent function which takes values on $\pi/2,\pi/2]$. This can be done by using the relationship $\text{atan}^{*}[\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{1/2}/\tau(\theta_{j},\theta_{k})]=\pi/2-\arctan[\tau(\theta_{j},\theta_{k})/\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{1/2}]$, which gives us an alternative expression for the density (18):
$c_{2}(\theta_{j},\theta_{k})=$ $\ \frac{1}{(2\pi)^{2}}\frac{1-\rho^{2}}{\{1-\tau(\theta_{j},\theta_{k})^{2}\}^{3/2}}$
$\times\left[\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}+\tau(\theta_{j},\theta_{k})\left\{\arctan\left(\frac{\tau(\theta_{j},\theta_{k})}{\sqrt{1-\tau(\theta_{j},\theta_{k})^{2}}}\right)-\frac{\pi}{2}\right\}\right].$
Since the density (18) is a function of $\theta_{1}-\operatorname{sign}(\rho_{12})\theta_{2}$, the density (18) is proportional to a univariate density on the circle, obtained by replacing $\tau(\theta_{j},\theta_{k})$ with $|\rho_{jk}|\cos\theta$, where $\theta\in[-\pi,\pi)$.
3.2 An extension of the copula case
The copula model (16) can be extended to have any univariate marginal distributions on the circle $\pi,\pi)$.
Definition 2.
Let a random vector $(U_{1},\ldots,U_{d})$ follow the proposed model $TPN_{d}(\bm{0},\bm{0},\bm{\Sigma})$. Let $f_{m,j}(\theta),\ j=1,\ldots,d,$ be densities on the circle with the mode at $\theta=0$ and $F_{m,j}(\theta)=\int_{0}^{\theta}f_{m,j}(t)dt$ be their cumulative distribution functions (CDFs). Define the circular random variables by
$\Theta_{1}=\mu_{1}+F_{m,1}^{-1}\left(\frac{U_{1}}{2\pi}\right),\quad\ldots,\quad\Theta_{d}=\mu_{d}+F_{m,d}^{-1}\left(\frac{U_{d}}{2\pi}\right).$ (19)
Then a new family of distributions on the hypertorus $\pi,\pi)^{d}$ is defined by the distribution of $(\Theta_{1},\ldots,\Theta_{d})$, which we call a copula-based extension of the (hyper-)toroidal projected normal (CTPN) distributions.
It immediately follows from Sklar’s theorem (e.g., Nelsen,, 2006, Theorem 2.10.9) that the random vector $(\Theta_{1},\ldots,\Theta_{d})$ defined in Definition 2 has the probability density function
$f(\bm{\theta})=(2\pi)^{d}\overline{c}(2\pi F_{m,1}(\theta_{1}-\mu_{1}),\ldots,2\pi F_{m,d}(\theta_{d}-\mu_{d}))\prod_{j=1}^{d}f_{m,j}(\theta_{j}-\mu_{j}),$ (20)
$\bm{\theta}=(\theta_{1},\ldots,\theta_{d})^{\top}\in[-\pi,\pi)^{d},$
where $\overline{c}$ denotes the density (16) with $\mu_{1}=\cdots=\mu_{d}=0$.
The equation (17) implies that the marginal density of $\Theta_{j}$ is
$f(\theta_{j})=f_{m,j}(\theta_{j}-\mu_{j}).$
The assumption that the mode of $f_{m,j}(\theta)$ is at $\theta=0$ ensures that the mode of each marginal of the CTPN distribution is given by $\theta_{j}=\mu_{j}$. The lower limit of the CDF $F_{m,j}$ is assumed to be zero because the mode of the joint density $f(\bm{\theta})$ of the CTPN distribution is located at $\bm{\theta}=(\mu_{1},\ldots,\mu_{d})$ if the copula density $\overline{c}(\bm{u})$ has the mode at $\bm{u}=\bm{0}$ (see Jones et al.,, 2015, Section 3.1). Flexible modeling is made possible through this extended model (20).
In addition it follows from Theorem 6 that the bivariate marginal distribution of $(\Theta_{j},\Theta_{k})$ has the density
$f(\theta_{j},\theta_{k})=(2\pi)^{2}\overline{c}_{2}(2\pi F_{m,j}(\theta_{j}),2\pi F_{m,k}(\theta_{k}))f_{m,j}(\theta_{j})f_{m,k}(\theta_{k}),$ (21)
where $\overline{c}_{2}$ denotes the density (18) with $\mu_{j}=\mu_{k}=0$. This bivariate case belongs to the general family of Wehrly and Johnson, (1980) with the circular density binding the two variables being $2\pi\overline{c}_{2}(\theta_{j},\theta_{k})$.
4 Coefficient of Correlation
(a)                            (b)
Figure 3: TNP distribution: plot of the circular correlation for different values of $\rho_{jk}$, and values of $\kappa_{j},\kappa_{k}\in\{0,0.49,2.45\}$: (a) $\kappa_{j}=\kappa_{k}$; and (b) $\kappa_{j}\neq\kappa_{k}$.
As seen in Figure 2, the parameter $\rho_{jk}$ governs the dependence between $\Theta_{j}$ and $\Theta_{k}$. To quantify how the value of $\rho_{jk}$ influences the dependence, we consider the circular correlation coefficient proposed by Rivest, (1982). This measure is considered here because it is a signed coefficient capable of distinguishing between positive and negative dependence. Let
$\xi_{jk}=\frac{\left|E(e^{i(\Theta_{j}-\Theta_{k})})\right|-\left|E(e^{i(\Theta_{j}+\Theta_{k})})\right|}{2},$
then the circular correlation coefficient between $\Theta_{j}$ and $\Theta_{k}$ proposed by Rivest, (1982) is defined as
$\rho_{jk,R}=\left\{\begin{array}[]{ll}\frac{\xi_{jk}}{\max\left\{E\left[\sin^{2}(\Theta_{j}-\mu_{j})\right],\ E\left[\sin^{2}(\Theta_{k}-\mu_{k})\right]\right\}},&E(e^{i(\Theta_{j}-\Theta_{k})})E(e^{i(\Theta_{j}+\Theta_{k})})\neq 0,\\ 2\xi_{jk},&E(e^{i(\Theta_{j}-\Theta_{k})})E(e^{i(\Theta_{j}+\Theta_{k})})=0.\end{array}\right$.
Figure 3 shows the circular correlation coefficient $\rho_{jk,R}$ for two cases: $\kappa_{j}=\kappa_{k}$ and $\kappa_{j}\neq\kappa_{k}$. The figure suggests that the circular correlation coefficient is a strictly increasing function of $\rho_{jk}$. The coefficient spans the entire range (1,1)(-1,1) when $\kappa_{j}=\kappa_{k}$, whereas its range is more limited when $\kappa_{j}\neq\kappa_{k}$. This can be inferred from the fact that if the range of $\rho_{jk}$ is extended and $\rho_{jk}=-1$ or $1$ in Definition 1, we have
$X_{cj}-\kappa_{j}=X_{ck}-\kappa_{k},\quad X_{sj}=\operatorname{sign}(\rho_{jk})X_{sk},$
and thus
$\Theta_{k}=\mu_{k}+\operatorname{sign}(\rho_{jk})\,\text{atan}^{*}\left(\frac{\sin(\Theta_{j}-\mu_{j})}{\cos(\Theta_{j}-\mu_{j})+(\kappa_{k}-\kappa_{j})/R_{j}}\right).$
This expression implies that $\Theta_{j}$ and $\Theta_{k}$ have perfect linear relationship when $\kappa_{j}=\kappa_{k}$. When $\kappa_{j}\neq\kappa_{k}$, the relationship between $\Theta_{j}$ and $\Theta_{k}$ depends on $R_{j}$, and the influence of $R_{j}$ increases as $|\kappa_{k}-\kappa_{j}|$ increases. This suggests the advantage of the CTPN distribution over the TPN distribution. Note, however, that the circular correlation coefficient $\rho_{jk,R}$ reaches 1-1 or $1$ when $\Theta_{j}$ and $\Theta_{k}$ have a linear relationship with probability one (see Rivest,, 1982, Proposition 2.3). Therefore, this coefficient is not suitable for measuring the relationship between the two variables when they exhibit a nonlinear dependence, which may occur when their marginal distributions differ.
In addition, it can be readily confirmed from Figure 3 that the positive and negative signs of $\rho_{jk}$ correspond to positive and negative values of the circular correlation coefficient, respectively.
5 MCMC Algorithm
In this section, we explain the algorithm that can be used to obtain posterior samples from the TPN distribution. To deal with the identification problem that the TPN inherited form the PN, we use the approach of Mastrantonio, (2018), which defines a model for the unidentifiable parameters and a posterior remap is used to obtain the identifiable version of the parameters.
While standard priors can be easily assigned to the parameters $\bm{\mu}$ and $\bm{\kappa}$, care must be taken when specifying a prior for $\bm{\Sigma}$ due to its relationship with $\bm{\Sigma}_{c}$. The Inverse Wishart (IW) distribution is a classical choice for modeling a covariance matrix, but not every valid covariance matrix $\bm{\Sigma}$ yields a valid $\bm{\Sigma}_{c}$. This makes its use problematic unless appropriate adjustments are made. The matrix $\bm{\Sigma}$ is defined over the domain $\mathcal{D}$ of all positive definite matrices, and a subset $\mathcal{D}^{*}$ of this domain can be defined as the set of all positive matrices that, when taking the absolute value of their elements, remain positive definite. We define a new distribution, which we call Truncated IW (TIW), as a distribution with value over $\mathcal{D}^{*}$ that has a density that is proportional to the IW. The normalization constant of this distribution is not known, but, as we will show, it is not needed in the MCMC algorithm.
The Bayesian model is as follows:
$\bm{\Theta}_{i}$ $\stackrel{{\scriptstyle iid}}{{\sim}}TPN_{d}(\bm{\mu},\bm{\kappa},\bm{\Sigma}),$ (22)
$\bm{\Sigma}$ $\sim TIW(\nu,\bm{\Psi}),\quad i=1,\dots,n,$ (23)
$\mu_{j}$ $\stackrel{{\scriptstyle iid}}{{\sim}}CU(0,2\pi),$ (24)
$\kappa_{j}$ $\stackrel{{\scriptstyle iid}}{{\sim}}N(\mathbf{m},\mathbf{v})I_{>0},\quad j=1,\dots,d,$ (25)
where $N(\mathbf{m},\mathbf{v})I_{>0}$ denotes a normal distribution with mean $\mathbf{m}$ and variance $\mathbf{v}$ truncated to the positive real line, while $CU(0,2\pi)$ is the uniform distribution on the circle $(0,2\pi)$. As we will show below, the truncated normal ensures that the full conditional of $\kappa_{j}$ is available in closed form, while for $\mu_{j}$, regardless of the distribution, no closed-form full conditional is available.
To make the inference easier, as is often done when working with the PN, the vectors of latent variables $\mathbf{R}_{i}=(R_{i,1},\dots,R_{i,d})^{\top}$ are introduced in the model. Then by defining the $j$-th element of $\mathbf{X}_{c}$ and $\mathbf{X}_{s}$ as
$[\mathbf{X}_{c}]_{j}$ $=R_{i,j}\cos(\Theta_{i,j}-\mu_{j}),\quad[\mathbf{X}_{s}]_{j}=R_{i,j}\sin(\Theta_{i,j}-\mu_{j}),$ (26)
we have that the joint density of $(\mathbf{R}_{i},\bm{\Theta}_{i})$ can be written as in (10), as
$f(\bm{r}_{i},\bm{\theta}_{i})=\frac{1}{(2\pi)^{d}|\bm{\Sigma}_{c}|^{1/2}|\bm{\Sigma}|^{1/2}}\exp\left[-\frac{1}{2}\left\{(\bm{x}_{c,i}-\bm{\kappa})^{\top}\bm{\Sigma}_{c}^{-1}(\bm{x}_{c,i}-\bm{\kappa})+\bm{x}_{s,i}^{\top}\bm{\Sigma}^{-1}\bm{x}_{s,i}\right\}\right]\prod_{j=1}^{d}r_{i,j}.$ (27)
5.1 MCMC steps for the TPN
For two of the steps of the MCMC algorithm, we need to use the conditional distribution of $x_{c,i,j}$ given all the other variables in the vector $\bm{x}_{c,i}$, which we indicate as $\bm{x}_{c,i,-j}$, and the conditional distribution of $x_{s,i,j}|\bm{x}_{s,i,-j}$. If we define $\bm{\Lambda}_{c}=\bm{\Sigma}_{c}^{-1}$, $\bm{\Lambda}_{s}=\bm{\Sigma}^{-1}$ and $\bm{\kappa}_{-j}$ as the vector $\bm{\kappa}$ without the $j$-the element, using standard results of a block matrix inversion and its connection with the conditional mean and variance of a multivariate normal, we know that
$x_{c,i,j}|\bm{x}_{c,i,-j}$ $\sim N(\eta_{c,i,j},\lambda_{c,j}^{-1}),\quad x_{s,i,j}|\bm{x}_{s,i,-j}\sim N(\eta_{s,i,j},\lambda_{s,j}^{-1}),$ (28)
$\eta_{c,i,j}$ $=\kappa_{j}-\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right),\quad\eta_{s,i,j}=-\lambda_{s,j}^{-1}\bm{\Lambda}_{s,j,-j}\bm{x}_{s,i,-j}$ (29)
where $\lambda_{c,j}$ and $\lambda_{s,j}$ are the $j$-th elements on the diagonal of $\bm{\Lambda}_{c}$ and $\bm{\Lambda}_{s}$ respectively, while $\bm{\Lambda}_{c,j,-j}$ and $\bm{\Lambda}_{s,j,-j}$ are the $j$-th rows of $\bm{\Lambda}_{c}$ and $\bm{\Lambda}_{s}$ respectively, without the $j$-th element.
Latent variable $r_{i,j}$.
We can easily see that the full conditional of $r_{i,j}$ is proportional to
$\exp\left[-\frac{(x_{c,i,j}-\eta_{c,i,j})^{2}}{2\lambda_{c,j}^{-1}}\right]\exp\left[-\frac{(x_{s,i,j}-\eta_{s,i,j})^{2}}{2\lambda_{s,j}^{-1}}\right]r_{i,j}$ (30)
which, by defining $\bm{\eta}_{i,j}=(\eta_{c,i,j},\eta_{s,i,j})^{\top}$, $\tilde{\bm{\Sigma}}_{i,j}=\text{diag}(\lambda_{c,j}^{-1},\lambda_{s,j}^{-1})^{\top}$, and $\mathbf{x}_{i,j}=(x_{c,i,j},x_{s,i,j})^{\top}$, can be written as
$\exp\left[-\frac{(\mathbf{x}_{i,j}-\bm{\eta}_{i,j})^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}(\mathbf{x}_{i,j}-\bm{\eta}_{i,j})}{2}\right]r_{i,j}$ (31)
which shows that $\bm{\theta}_{i,j}|\dots\sim PN(\bm{\eta}_{i,j},\tilde{\bm{\Sigma}}_{i,j})$, and hence we can use the algorithm proposed by Hernandez-Stumpfhauser et al., (2017). First, we define
$\bm{\xi}_{i,j}=(\cos(\theta_{i,j}),\sin(\theta_{i,j}))^{\top},\quad A_{i,j}=\mathbf{u}_{i}^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}\bm{\xi}_{i,j},\quad B_{i,j}=\bm{\xi}_{i,j}^{\top}\tilde{\bm{\Sigma}}_{i,j}^{-1}\bm{\eta}_{i,j}.$
Then, we draw ${\beta}_{1,i,j}\sim U\left(0,\exp\left(-\frac{1}{2}A_{i,j}\left(r_{i,j}-\frac{B_{i,j}}{A_{i,j}}\right)^{2}\right)\right)$ and ${\beta}_{2,i,j}\sim U(0,1)$, and the new sample of $r_{i,j}$ is
$r_{i,j}=\left((\varrho_{2,i,j}^{2}-\varrho_{1,i,j}^{2}){\beta}_{2,i,j}+\varrho_{1,i,j}^{2}\right)^{1/2},$
where
$\varrho_{1,i,j}=\frac{B_{i,j}}{A_{i,j}}+\max\left\{-\frac{B_{i,j}}{A_{i,j}},-\sqrt{\frac{-2\log{\beta}_{1,i,j}}{A_{i,j}}}\right\},\quad\varrho_{2,i,j}=\frac{B_{i,j}}{A_{i,j}}+\sqrt{\frac{-2\log{\beta}_{1,i,j}}{A_{i,j}}}.$
Parameter $\mu_{j}$.
Being $\mu_{j}$ embedded inside the sine and cosine used to compute $\mathbf{X}_{c}$ and $\mathbf{X}_{s}$, it is not generally possible to find a closed form expression for its full conditional, which must be then sampled using a Metropolis Hastings algorithm.
Parameter $\kappa_{j}$.
Since the kernel of the prior distribution of $\kappa_{j}$ is the same as the one of a normal distribution, from the conditional distribution (28) we can see that
$x_{c,i,j}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right)\sim N(\kappa_{j},\lambda_{c,j}^{-1}).$
Hence, the full conditional of $\kappa_{j}$ is proportional to
$\exp\left[-\frac{1}{2}\left(\kappa_{j}^{2}v^{-1}-2\kappa_{j}v^{-1}m+\kappa_{j}^{2}n\lambda_{c,j}-2\kappa_{j}\sum_{i=1}^{n}\left(\bm{x}_{c,i}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}\left(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j}\right)\right)\right)\right]$ (32)
$\propto\exp\left[-\frac{(\kappa_{j}-m_{p,j})^{2}}{2v_{p,j}}\right],$ (33)
where
$v_{p,j}$ $=(v^{-1}+n\lambda_{c,j})^{-1},$
$m_{p,j}$ $=v_{p,j}\left(\sum_{i=1}^{n}\left(\bm{x}_{c,i}+\lambda_{c,j}^{-1}\bm{\Lambda}_{c,j,-j}(\bm{x}_{c,i,-j}-\bm{\kappa}_{-j})\right)+v^{-1}m\right).$
It is easy to see that this is the kernel of $N(m_{p,j},v_{p,j})$, and since $\kappa_{j}>0$, the full conditional of $\kappa_{j}$ is $N(m_{p,j},v_{p,j})\mathbf{I}_{\kappa_{j}>0}$.
Parameter $\bm{\Sigma}$.
To sample this parameter, we use a Metropolis–Hastings step. At iteration $b$, the proposal distribution is set to $IW(\tau,\bm{\Sigma}^{b}(\tau-(d+1)))$, where $\bm{\Sigma}^{b}$ denotes the most recently accepted value of the parameter in the MCMC. The expected value of this proposal distribution is $\bm{\Sigma}^{b}$, and as the parameter $\tau$ increases, the variance of the distribution decreases, making $\tau$ a tuning parameter that controls the variability of the proposal.
After $\bm{\Sigma}^{*}$ is proposed, we compute $\bm{\Sigma}_{c}^{*}$ and if this is not a positive definite matrix, the proposal is rejected and we move to the next step of the algorithm. Otherwise, we use the standard Metropolis machinery to decide whether to accept or reject the sample.
5.2 General copula approach
Let $c$ be the density of the $d$-variate TPN with uniform margins ($TPN(\mathbf{0},\mathbf{0},\bm{\Sigma})$). If we want to use the proposed density as a copula model, with $j$-th marginal density equal to $f_{m,j}(\theta_{i,j})$ and CDF $F_{m,j}(\theta_{i,j})$, we have that the joint density of observation $i$ is equal to (20). Since $c$ has the functional form of a $TPN_{d}(\mathbf{0},\mathbf{0},\bm{\Sigma})$, it can be seen as the result of the integral
$c\left(\theta_{i,1}^{*},\dots,\theta_{i,d}^{*}\right)=\int_{(\mathbb{R}^{+})^{d}}f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})\prod_{j=1}^{d}r_{i,j},$ (34)
where $f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})$ and $f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})$ in this case are the density of a $d$-variate normal with mean vector of zeros, covariance matrix $\bm{\Sigma}_{c}$ and $\bm{\Sigma}$ respectively, and
$\theta_{i,j}^{*}$ $=2\pi F_{m,j}(\theta_{i,j}-\mu_{j}),$ (35)
$\bm{x}_{c,i}^{*}$ $=\left(r_{i,1}\cos(\theta_{i,1}^{*}),\dots,r_{i,d}\cos(\theta_{i,d}^{*})\right),$ (36)
$\bm{x}_{s,i}^{*}$ $=\left(r_{i,1}\sin(\theta_{i,1}^{*}),\dots,r_{i,d}\sin(\theta_{i,d}^{*})\right).$ (37)
This means that if we introduce the latent variables $\mathbf{r}_{i}=(r_{i,1},\dots,r_{i,d})^{\top}$, we can work with the joint density
$f(\bm{\theta}_{i},\mathbf{r}_{i})=\left(f_{\bm{X}_{c,i}^{*}}(\bm{x}_{c,i}^{*})f_{\bm{X}_{s,i}^{*}}(\bm{x}_{s,i}^{*})\prod_{j=1}^{d}r_{i,j}\right)(2\pi)^{d}\prod_{j=1}^{d}f_{m,j}(\theta_{i,j}-\mu_{j}).$ (38)
Hence, the parameter $\bm{\Sigma}$ and the latent variables $\mathbf{r}_{i}$ can be sampled as in the previous section, whereas the variables $\bm{\theta}_{i}^{*}$, $\bm{x}_{c,i}^{*}$, and $\bm{x}_{s,i}^{*}$ must be used in place of $\bm{\theta}_{i}$, $\bm{x}_{c,i}$, and $\bm{x}_{s,i}$. The marginal density can depend on some parameters, that must be sampled using the Metropolis-Hastings algorithm, if closed-form full conditionals are not available.
6 Real Data Example
Figure 4: Map showing the locations of four buoys in the Gulf of Mexico. A grid with 1-degree intervals of latitude and longitude is overlaid on the map to provide geographic reference.
Figure 5: Scatterplot of all possible combinations of observed data.
Figure 6: Posterior density of parameter $\mu_{k,d}$ for the 6 circular variables. The y-axis is represented in logarithmic scale to better visualize the densities.
Figure 7: Posterior density of parameter $\lambda_{k,d}$ for the 6 circular variables. The y-axis is represented in logarithmic scale to better visualize the densitiess.
Figure 8: Heatmap of the posterior means of matrix $\bm{\Sigma}$.
To demonstrate the practical applicability of the proposed distribution, in this section, we apply it to a real dataset obtained from the National Data Buoy Center website (http://www.ndbc.noaa.gov/). We use data from four buoys located in the Gulf of Mexico, an ocean basin and a marginal sea of the Atlantic Ocean, off the Texas coast. Two buoys (Station PTAT2: 27.826° N, 97.051° W, and Station 42035: 29.235° N, 94.410° W) provide data only on wind direction, while the other two (Station 42019: 27.908° N, 95.343° W, and Station 42020: 26.970° N, 96.679° W) include both wind and wave directions: the spatial location of the buoys is shown in Figure 4. The data we use are daily measurements taken at 6 PM, from January 1st, 2020 to December 31st, 2020, yielding 366 recordings of a 6-dimensional vector of circular random variables. In the selected time window, station PTAT2 has 1 missing value. Station 42035 has 92 missing values, station 42019 has 157 missing values for wind direction and 57 for wave direction, while station 42020 has 46 missing values in wave direction and 51 in wind direction. The missing values are treated as additional random variables, and posterior values are simulated during the MCMC algorithm. A scatterplot for each pair of data is shown in Figure 5.
By using the proposed MCMC, we estimate the parameters of a 6 dimensional TPN as well as a copula model with wrapped Cauchy (WC) marginals. The WC density is formalized as
$f(\theta_{i,j}|\mu_{j}^{*},\lambda_{j})=\frac{1}{2\pi}\cdot\frac{1-\lambda_{j}^{2}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(\theta-\mu_{j}^{*})}$ (39)
and it depends on the mean direction $\mu_{j}^{*}$, which is also the mode of the distribution, and a concentration parameter $\lambda_{j}\in[0,1]$. Its CDF can also be written in closed form as
$F(\theta_{i,j}\leq a|\mu_{j}^{*},\lambda_{j})=\begin{cases}\frac{1}{2\pi}\arccos\left(\frac{(1+\lambda_{j}^{2})\cos(a-\mu_{j}^{*})-2\lambda_{j}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(a-\mu_{j}^{*})}\right),&\text{if }\sin(a-\mu_{j}^{*})\geq 0\\ 1-\frac{1}{2\pi}\arccos\left(\frac{(1+\lambda_{j}^{2})\cos(a-\mu_{j}^{*})-2\lambda_{j}}{1+\lambda_{j}^{2}-2\lambda_{j}\cos(a-\mu_{j})}\right),&\text{if }\sin(a-\mu_{j}^{*})<0\end{cases}$ (40)
where $a\in[0,2\pi)$. It should be notated that, by construction, the CDF at $\mu_{j}^{*}$ is equal to zero. Interestingly, it is possible to find in closed form also the quantile function of the WC which is
$F^{-1}(u|\mu_{j}^{*},\lambda_{j})=\left(\mu_{j}^{*}+\operatorname{sign}\left(0.5-u\right)\cdot\arccos\left(\frac{2\lambda_{j}+(1+\lambda_{j}^{2})\cos(2\pi u)}{1+\lambda_{j}^{2}+2\lambda_{j}\cos(2\pi u)}\right)\right).$ (41)
Equations (40) and (41), under the assumption $\mu_{j}^{*}=0$, can be used in (19). Since the WC is symmetrical, $\mu_{j}$ is the circular mean of the $j$-th marginal distribution.
We assume circular uniform prior distributions for the mean directions of both models, we assume $\kappa_{j}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,100000)I_{>0}$, $\lambda_{j}\sim U(0,1)$ and $\bm{\Sigma}\sim TIW(8,\mathbf{I}_{6})$. The MCMC is implemented with 30,000 iterations, with a burn-in of 10,000 and a thinning factor of 10. To select the best model between the two, we use the Continuous Ranked Probability Score (CRPS) (Gneiting and Raftery,, 2007). The CRPS is a metric that compares the true value of a random variable with its predictive distribution and is straightforward to compute in a Bayesian setting. In this real data example, it is computed using 10% of the data, which is set aside and not used in the model fitting. The CRPS for the TPN is 0.133 while for the WC is 0.117: since the WC is the one with lower CRPS index, results are only presented for this model.
In terms of both the mean direction $\mu_{k,d}$ (Figure 6) and the concentration parameter $\lambda_{k,d}$ (Figure 7), the distributions of the two wave variables appear quite similar, which is not the case for the wind directions. Figure 8 highlights a strong correlation between variables of the same type, while the correlation between wave and wind directions is weaker, though still around 0.5 for all combinations, even when recorded at the same buoy.
7 Conclusion
In this paper, we introduced two families of probability distributions designed for the Bayesian analysis of hypertoroidal data. The first family, the TPN distribution, is derived through the projection of multivariate normal distributions under specific parameter constraints. This construction ensures that the TPN distribution maintains key tractable properties, such as closure under marginalization and an intuitive parameter interpretation. The second family is the CTPN model, a copula-based extension of the TPN model, offering additional flexibility by allowing arbitrary univariate marginal distributions. Our proposal is the first copula model for vectors of any dimension of circular data using the projection approach.
The proposed distributions address some of the limitations of existing models, particularly by providing a clear interpretation of parameters governing the mean directions, concentration, and dependence structure. The use of the Markov Chain Monte Carlo (MCMC) framework for Bayesian estimation further enhances the applicability of these models, as demonstrated through simulations and real data applications.
Acknowledgments
This work was partially supported by Mitsui Sumitomo Insurance Co., Ltd. (to Shogo Kato and Gianluca Mastrantonio) and JSPS KAKENHI Grant Number 25K15037 (to Shogo Kato).